OpenStudy (anonymous):

Convert the mixed number to an improper fraction. 4 7/8

OpenStudy (lieutenantgeneral):

4 7/8 = 8 * 4 = 32 + 7 = 39 so 39/8

OpenStudy (anonymous):

It equals 39/8